# Mapping the Global Election Landscape on Social Media in 2024

### v1

@@ Introduction
We implemented a two-step process to assess the topics discussed in each post. We processed each country separately since each one had posts in their primary language. For each country, we selected a random sample of 50,000 posts made by news organizations. We then cleaned the text of these posts by removing links, hashtags, and non-alphanumeric characters, such as emojis. Subsequently, we applied the BERTopic clustering algorithm to perform topic clustering for each country. The resulting clusters were labelled with overarching themes such as ‘economy,’ ‘crime,’ ‘entertainment,’ and ongoing conflicts in Ukraine and the Middle East (identified as ‘mena’ in the analysis). Notice that, Finland and Sweden did not produce any discernible topics during the clustering phase. To infer the topic of all posts, we computed the term frequency-inverse document frequency (tf-idf) values of the words from the posts within each cluster, treating the concatenation of all posts from each macro-topic as a single document. This allowed us to identify the macro-topic that most closely aligned with the textual content of each post based on the tf-idf scores of the words. Fig. 1 presents the share of posts discussing each topic. As shown, political parties predominantly post about internal politics, whereas ‘lighter’ topics such as entertainment or sports garner significantly more attention from news outlets.
@@
Interestingly, the number of reactions received by posts from political parties and news agencies falls within similar orders of magnitude, with tails ranging between $10^{3}$ and $10^{5}$ reactions. Engagement distribution patterns for each country are displayed in Fig. S3 for news agencies and S4 for political parties, exhibiting the same consistent patterns.
@@
We focus on the news outlets case (the parties are treated similarly). Following a method similar to the previous one, we consider a country $c$ with $n_{c}$ labeled news outlets discussing $m_{c}$ topics. For each country, we construct a matrix $M^{c}$ of dimension $n_{c}\times m_{c}$, where $M^{c}_{ij}$ represents the fraction of engagement received on topic $j$ by news outlet $i$. Fig. 5 shows the biplot resulting after applying PCA to $M^{c}$ on each country having $n_{c}\geq 3$.
@@
Interestingly, the analysis reveals that for both news outlets and political parties, the political leaning of each entity does not significantly influence how engagement is distributed across topics. This suggests that no topics are exclusively engaged by entities with a specific political leaning.
One possible explanation for the observed engagement patterns across different political leanings can be attributed to the overall narrative style of each page. Even when discussing the same topics, different pages may present perspectives that resonate more closely with specific audiences. However, due to the limitations of our classification approach, PCA primarily captures the general content themes without distinguishing between nuanced perspectives.
Despite this limitation and the potential imbalances in labeling across different countries, our findings suggest that news outlets and political parties, regardless of their political orientation —left, center, or right— are likely to engage with similar topics.
@@
We observe that, for what concerns left news outlets Poland, Spain, USA, Ireland and UK are getting high engagements in topic such as abort, religion, climate and tech. At the same time Netherlands, France and Germany are more focused on education, economy, africa and migration. On the other hand, right news outlet in Italy, Germany, UK and Netherlands are getting high engagements in topics such as gender, migration and ukraine. Interestingly, USA, Portugal, Denmark and Spain are more interested in religion, abortion, sport and crime. Finally, only right news outlets from France and Austria seems to get engagement with the climate topic.
We observe a more heterogeneous scenario for political parties, where all topics are treated more homogeneously.
@@ Acknowledgments
Fig S3: Reaction distributions to news posts across countries.

### v2

@@ Introduction
We implemented a two-step process to assess the topics discussed in each post. We processed each country separately since each one had posts in their primary language. For each country, we selected a random sample of 50,000 posts made by news organizations. We then cleaned the text of these posts by removing links, hashtags, and non-alphanumeric characters, such as emojis. Subsequently, we applied the BERTopic clustering algorithm to perform topic clustering for each country. The resulting clusters were labelled with overarching themes such as ‘economy,’ ‘crime,’ ‘entertainment,’ and ongoing conflicts in Ukraine and the Middle East (identified as ‘mena’ in the analysis). Notice that, Finland and Sweden did not produce any discernible topics during the clustering phase. To infer the topic of all posts, we computed the term frequency-inverse document frequency (tf-idf) values of the words from the posts within each cluster, treating the concatenation of all posts from each macro-topic as a single document. This allowed us to identify the macro-topic that most closely aligned with the textual content of each post based on the tf-idf scores of the words.
Fig. 1 presents the share of posts discussing each topic. As shown, political parties predominantly post about internal politics, whereas ‘lighter’ topics such as entertainment or sports garner significantly more attention from news outlets.
@@
Interestingly, the number of reactions received by posts from political parties and news agencies falls within similar orders of magnitude, with tails ranging between $10^{3}$ and $10^{5}$ reactions. Engagement distribution patterns for each country are displayed in Fig. S1 for news agencies and S2 for political parties, exhibiting the same consistent patterns.
@@
We focus on the news outlets case (the parties are treated similarly). Following a method similar to the previous one, we consider a country $c$ with $n_{c}$ labeled news outlets discussing $m_{c}$ topics. We construct the matrix $M^{c}$ of dimension $n_{c}\times m_{c}$, where $M^{c}_{ij}$ represents the fraction of engagement received on topic $j$ by news outlet $i$. Fig. 5 shows the biplot resulting after applying PCA to $M^{c}$ on each country having $n_{c}\geq 3$.
@@
Interestingly, the analysis reveals that the political leaning of each entity does not significantly influence how engagement is distributed across topics.
One possible explanation for the observed engagement patterns can be attributed to the overall narrative style of each page. Even when discussing the same topics, different pages may present perspectives that resonate more closely with specific audiences. However, due to the limitations of our classification approach, PCA primarily captures the general content themes without distinguishing between nuanced perspectives.
Despite this limitation and the potential imbalances in labeling across different countries, our findings suggest that news outlets and political parties with also different political leaning —left, center, or right— are likely to engage with similar topics.
@@
We observe that left news outlets in Poland, Spain, USA, Ireland and UK are getting high engagement in topics such as abortion, religion, climate and tech. At the same time Netherlands, France and Germany are more focused on education, economy, africa and migration. On the other hand, right news outlet in Italy, Germany, UK and Netherlands are getting high engagement in topics such as gender, migration and ukraine. Interestingly, USA, Portugal, Denmark and Spain are more interested in religion, abortion, sport and crime. Finally, only right news outlets from France and Austria seem to get engagement with the climate topic.
Fig. 6 $(c)-(d)$ depict the results of the same analysis for Political Parties. In this case we observe even more differences between each country, suggesting that political parties from different countries, even if with the same political leaning - left or right - tend to engage in varying and different topics.
Finally, we are interested in highlighting the topics discussed by each political leaning in each country. To hit this aim, we employ the following procedure: For each country $c$, we construct a matrix $M^{c}$ of dimension $3\times m_{c}$ where the rows represent the political leanings of the news outlets—left, center, and right. The entry $M^{c}_{ij}$ denotes the fraction of engagement received by news outlets with political leaning $i$ in topic $j$. PCA is then applied to this matrix to yield a two-dimensional representation capturing the most significant aspects of the data.
Fig 7: Most engaging topics of news outlets by political leaning.
The resulting biplots generally indicate that right-leaning news outlets attract more engagement on topics related to politics, religion, and migration, whereas left-leaning news outlets are more actively engaged in topics like education, health, and technology. Additionally, the angles between topics suggest a slight negative correlation between the topics favored by the left and right, indicating a moderate level of, overall, topic segregation.
Similarly, Fig. 8 displays the engagement patterns by political leaning for a subset of countries in the parties dataset. The complete results are detailed in Fig. S4.
Fig 8: Most engaging topics of parties by political leaning.
@@ Acknowledgments
The work is supported by IRIS Infodemic Coalition (UK government, grant no. SCH-00001-3391), SERICS (PE00000014) under the NRRP MUR program funded by the European Union - NextGenerationEU, project CRESP from the Italian Ministry of Health under the program CCM 2022, PON project “Ricerca e Innovazione” 2014-2020, and PRIN Project MUSMA for Italian Ministry of University and Research (MUR) through the PRIN 2022. This work was supported by the PRIN 2022 “MUSMA” - CUP G53D23002930006 - Funded by EU - Next-Generation EU – M4 C2 I1.1.
@@
Fig S1: Reaction distributions to news posts across countries.